OpenStudy (anonymous):

Find two positive numbers whose product is 4 and such that 4x+2y is a minimum. a)2 and 2 b)(1/4) and 4 c) (1/2) and 8 d) sqrt of 3 and (4/sqrt of 3) e) sqrt of 2 and (2*sqrt of 2)

OpenStudy (anonymous):

The answer is e) \[ y = \frac 4 x\\ f(x) = 4 x + 2y= 4 x + \frac 8 x\\ f'(x)= 4-\frac{8}{x^2}\\ f'(x)=0 \text { for} x =\pm \sqrt 2\\ f''( \sqrt 2) > 0\\ \] The minimum is at \[ x= \sqrt 2\] and y will be \[ y= \frac 4 {x}= 2 \sqrt 2\]